# Linearly-scalable and entropy-optimal learning of nonstationary and nonlinear manifolds

### v2

@@ 2.3 Incorporating simultaneous learning of data reliability in EOMC
Learning piece-wise low-dimensional linear manifold approximations with EOMC implicitly deploys an assumption that all of the data instances have the same uniform probability $(1/T)$ to contribute to the loss function $\mathcal{L}^{\textrm{EOMC}}$ in (2.2). This is not a problem when this assumption is justified, and when the training data is spread more-or-less uniformly and close to the manifold. But, this becomes a problem when the new data instances $X$ (f.e., the ones that were not used in the training), come from somewhere arbitrarily in the feature space - and further away from the manifold. In the Sec.3.1.3 below, we provide an example of such a situation - resulting in projection artefacts, when training the manifold with the data close to the manifold (red crosses in Fig. 4), and then using the trained EOMC manifold projectors (11) from Theorem 1 for the data points that are far away from the manifold and from the training data (blue dots in Fig. 4B). This input data reliability issue is a very general and severe problem in contemporary AI, resulting in low robustness, "over-confidence" in provided answers, and in AI vulnerability to the so-called adversarial attacks, when subject to new input data that is sufficiently-different from the domain $\Omega$ of "admissible" data implicitly defined by the training data.
To address this issue - and following the ideas originally introduced in the Entropic Outlier Sparsification (EOS) and Entropy-Optimal Networks (EON) methods [21, 24] - we propose to incorporate in the EOMC problem formulation (2.2) an additional learnable variable $\gamma_{0}(t)=\mathbb{P}\left[X(:,t)\in\Omega\right]$, measuring a probability that a new data instance $X(:,t)$ belongs to the "admissible" data domain $\Omega$. Denoting $l(t)=-\sum_{k=1}^{K}\gamma(k,t)\left[\left(X(:,t)-\mu_{k}\right)^{\dagger}\left(\left(I+\alpha\right)-\mathcal{T}_{k}\mathcal{T}_{k}^{\dagger}\right)\left(X(:,t)-\mu_{k}\right)+\beta\textrm{log}_{K}\left(\gamma(k,t)\right)\right]$, we can now rewrite (2.2) as a maximisation of the expectation $\mathbb{E}_{\gamma_{0}}\left[l\right]$ of the loss $l$ wrt. this probability measure $\gamma_{0}$, subject to the Shannon entropy regularization terms $\beta_{0}\gamma_{0}\left(t\right)\textrm{log}_{T}\gamma_{0}\left(t\right)$, where $\beta_{0}\geq 0$. Re-arranging terms we obtain:
$\left\{\gamma_{0}^{*},\gamma^{*},\mu^{*}_{1},\mathcal{T}^{*}_{1},\dots,\mu^{*}_{K},\mathcal{T}^{*}_{K}\right\}=\arg\min\mathcal{L}_{\gamma_{0}}^{\textrm{EOMC}}$
$\mathcal{L}_{\gamma_{0}}^{\textrm{EOMC}}$ $=$ $\sum_{t=1}^{T}\gamma_{0}(t)\left[\beta_{0}\textrm{log}_{T}\gamma_{0}\left(t\right)-l(t)\right],$
@@
Shannon entropy regularization terms $\beta_{0}\gamma_{0}\left(t\right)\textrm{log}_{T}\gamma_{0}\left(t\right)$ are incorporated here with a positive sign, since for $\beta_{0}\equiv 0$, $\gamma_{0}$ solution of (2.3) is entropy-minimal, attaining only $1/0$ values - and the entropy-optimal $\gamma_{0}$-solutions can only be found when maximizing the Shannon entropy. It is also straightforward to validate that the previous formulation (2.2) becomes a special case of (2.2) when $\beta_{0}\rightarrow+\infty$ (since then $\gamma_{0}(t)\rightarrow(1/T)$, for all $t$).
Then, following the same line of argumentation as in the proof of Theorem 1 above, it is easy to show that for the fixed values of the loss $l$ (hence, when all other learning variables except of $\gamma_{0}$ are fixed), and for any $\beta_{0}>0$, there exists a unique solution $\gamma^{*}_{0}$ of the problem (2.3):
@@ 2.4 Selection of EOMC hyper-parameters $d,K,\alpha,\beta,\beta_{0}$
As will be demonstrated below on practical examples (for example, see Fig. 1), tuning the EOMC hyper-parameters $d$ and $K$ can be done iteratively. To select the optimal values for $d$ and $K$, for each of the $k=1,\dots,K$ manifolds one inspects the decay of eigenvalues $\Lambda^{(k)}_{1,1},\Lambda^{(k)}_{2,2},\dots,\Lambda^{(k)}_{D,D}$ for the local weighted covariances $\mathrm{Cov}_{k}\left(X,\gamma,\mu_{k}^{*}\right)$ in the local manifolds: as follows from the Theorem 1 above, $\sum_{i=d+1}^{D}\Lambda^{(k)}_{i,i}$ quantifies the amount of squared 2-norm loss from projecting the data, that is "assigned" to this manifold $k$, through the EOMC internal coordinates $\gamma(k,:)$. Hence, it quantifies the local quality of lossy compression, that is achieved when projecting the high-dimensional data on the local low-dimensional manifolds. Since (2.2) is an unsupervised learning problem, we can not apply most of the standard hyper-paremeter selection routines from ML and AI, like cross-validation and Bayesian hyper-paremeter tuning [25]. Instead, one can use here the hyper-parameter selection from the unsupervised regularization methods in statistics and computational science, like the $L$-curve method [26]. Alternatively, in the following examples we will adopt a nonparametric information-theoretic perspective to model selection - aiming to find the hyper-parameter combinations that lead to the models combining simplicity (measured as a high lossy compression rate, computed as a ratio between the raw data complexity and EOMC model descriptor length), and quality (measured as the relative loss of compression). In contrast to common parametric measures from information theory (like Akaike and Bayesian Information Criteria) [27] that rely on validity of parametric assumptions like Gaussianity, this non-parametric model selection procedure based on lossy compression allows a robust identification of good hyper-parameter combinations across all of the benchmarks considered below. As will also be shown on application examples below, EOMC does not require a careful and precise adjustment, since the results remain robust in the broad ranges of hyper-parameters.
@@ 3.1 Synthetic examples
  • Phase 1: select the hyper-parameters $d,K,\alpha,\beta$ (and $\beta_{0}$, if simultaneous learning of data reliability $\gamma_{0}$ with (2.3) is incorporated in EOMC [^3] Here, and in the next example described in Sec. 3.1.2, we perform the analysis setting $\gamma_{0}\equiv 1/T$ and $\beta_{0}=0$. An example of EOMC with simultaneous learning of the entropy-optimal data reliability function $\gamma_{0}$ will be provided in Sec. 3.1.3), as described in Sec. 2.4. Next, perform the EOMC analysis according to the Theorem 1 above. Note that the application of the standard linear PCA (left panel in Fig. 1) does not reveal any spectral gap in the eigenvalues of the covariance, and does not indicate a presence of any low-dimensional manifold. In contrast, applying EOMC with arbitrarily selected $K=15$ indicates a clear spectral gap after the second eigenvalues of the localized EOMC covariance matrices $\mathrm{Cov}_{k}\left(X,\gamma,\mu_{k}^{*}\right)$. This clearly indicates that the data contains non-linear manifolds with local dimensionality $d=2$.
@@
Despite of the hyper-parameter adjustment, PCA and t-SNE fail to recover the two low-dimensional manifolds from these nonstationary data (see Fig. 2A and 2B). Fig. 2C illustrates further effects of hyper-parameter selection described in Sec. 2.4: setting both $\alpha=0$ and $\beta=0$ makes (2.2) to its special case, i.e., to PCA-clustering [16, 17, 18]. It makes visible the central problem of PCA-clustering, discussed above and induced by the Lemmas 1 and 2: the kernel of the manifold projection operator is non-empty when $d<D$ (Lemma 1), and the errors orthogonal to the manifold are not visible to the method. Therefore, the approximation in this case is rather cutting through the manifold then approximating it. Setting $\alpha$ and $\beta$ to some non-zero values mitigates this problem (see Fig. 2D). Then, increasing the number $K$ of local manifolds from 15 (which was the initial guess) to 50 results in almost perfect reconstruction of both nonlinear manifolds (see Fig. 2D). Please note that obtaining this result did not require a tedious hyper-parameter tuning, results shown in Fig. 2C were obtained from only two repetitions of the EOMC pipeline described in the Fig. 1.
@@
As in the previous example, despite of the hyper-parameter tuning, both linear PCA and t-SNE fail to find these low dimensional manifolds hidden in 100 dimensions of noisy data (see Fig. 3A and 3B). Instead, t-SNE finds a lot of clusters (see Fig. 3B), that are not present in the data, representing a clear artefact of t-SNE. Fig. 3C illustrates again the problem of PCA-clustering, and Figs. 3D and 3E show the effect of the entropic regularization term in (2.2): setting $\alpha>0$ and $\beta=0$ results in "edgy" piece-wise linear approximation of the nonlinear manifold fragments (Fig. 3D), whereas setting $\alpha>0$ and $\beta>0$ allows obtaining much better and more smooth interpolations. Like in the example 1 above, obtaining the result from the Fig. 3E did not require a tedious hyper-parameter adjustment, requiring to go through the EOMC pipeline described in the Fig. 1 only once.
@@
Next, we illustrate the application of the modified EOMC (with a simultaneous learning the input data reliability $\gamma_{0}$), as described in Sec.  2.3 and eqs. (2.3-15). We generate a noisy 2D data (red crosses in Fig. 4A), that are normally distributed around a smooth 1D non-linear manifold following a smooth letter "S" (black line in Fig. 4A). As in the previous synthetic examples, to make the manifold learning more challenging, we add an additional 10-dimensional Gaussian noise and randomly rotate the training data in 10 dimensions. As discussed in Sec.  2.3, the EOMC formulation in (2.2) implicitly assumes that all of the training data points have the same $1/T$ contribution to the loss function - i.e., according to the uniform probability distribution prior. However, this assumption can be seriously violated when considering the data points not used in training and distributed uniformly inside of the whole 10-dimensional hypercube - and far away from the training data (blue dots in Fig. 4). As can be seen from Fig. 4B, applying the projection formula (11) from the Theorem 1 results in artefact projections that are far away of the true manifold (red circular dots in Fig. 4B).
In contrast, co-inference of data reliability function $\gamma_{0}(t)=\mathbb{P}\left[X(:,t)\in\Omega\right]$ in the modified EOMC from Sec.  2.3 and eqs. (2.3-15) produces a very reliable reconstruction of the true manifold: the intensity of red for circular dots in Fig. 4C is proportional to the inferred values of data reliability $\gamma_{0}(t)$. The unreliable points far away from the training data and from the EOMC-approximated manifold automatically get values of $\gamma_{0}(t)$ close to zero. As in the previous two synthetic examples, achieving this result did not require any particular hyper-parameter tuning: it was obtained for an ad hoc choice of $K=6,d=1,\alpha=5,\beta=0.3,\beta_{0}=4$. This result remains remarkably-robust, and does not change notably when changing these hyper-parameter values in quite broad ranges.
Figure 4: Illustration of the modified EOMC learning from Sec.  2.3 and eqs. (2.3-15): co-inference of data reliability function $\gamma_{0}$.
@@ 3.2 Analysis of data from Lorenz-63 in chaotic, strongly-chaotic and very strongly-chaotic regimes
We inspect the time series of intrinsic EONC coordinates $\gamma$ that were computed with these optimal hyper-parameter choices. $\gamma$ encode the probabilities of points to belong to different local manifolds $\left\{\mu_{k},\mathcal{T}_{k}\right\}$, $k=1,\dots,K$ at different time series instances. As can be seen from the right column of Fig. 5, $\gamma$ exhibits very persistent and metastable dynamics in time, indicating relatively-long stays in each of the local linear manifolds (with $d=3$). Please note that this metastability of $\gamma$ can not be an artefact of the EOMC data analysis: as can be seen from (2.2), the EOMC loss function does not contain any terms that would enforce the persistence or metastability on $\gamma$. As a matter of fact, the value of $\mathcal{L}^{\textrm{EOMC}}$ in (2.2) is invariant with respect to any permutation of the columns of data matrix $X$. This means that the observed metastability of $\gamma$ can only be an imprint of the underlying L96 dynamics - that appears to be best described by a metastable process switching between low-dimensional (with $d=3$) locally-linear manifolds. Next, for each of the obtained $\gamma$, we compute the Markovian transition operators $P$ that describe the time evolution of $\gamma(:,t)$, by means of the exact law of the total probability, i.e., $\gamma(:,t+\tau)=P\gamma(:,t)$, where $P_{i,j}=\mathbb{P}\left[\gamma(j,t+\tau)=1|\gamma(i,t)=1\right]$ [34]. Elements $P_{i,j}$ of the transfer operator $P$ contain the probabilities of transitions from state $j$ to state $i$ in a single time step $\tau$. As demonstrated by the Fig. 6A, operators $P$ are characterized by high probabilities of staying in the states (large diagonal entries) and low probabilities of transitions to other local manifold states (very small off-diagonal entries). Next, for each of the three transfer operators computed for each of the three L96 forcing regimes, we compute the relaxation times $\tau^{rel}(P)=\frac{\tau}{1-|\lambda_{2}|}$, where $\tau$ is the time step, and $\lambda_{2}$ is the second largest (in absolute value) eigenvalue of $P$ [34]. Relaxation times $\tau^{rel}(P)$ in Markov processes measure the predictability horizons: they quantify the time it takes for the Markov processes $P$ to forget the initial condition, and to converge towards the invariant density measure [35, 36, 34]. As can be seen from the Fig. 6A, $\tau^{rel}(P)$ gradually reduces from $0.73$, over $0.61$ and to $0.57$ when $F$ goes from $7$, over $9$, and to $12$.
@@
$\frac{\partial\zeta}{\partial t}+\{\phi,\zeta\}$ $=\alpha(\tilde{\phi}-\tilde{n})-\mu\nabla^{4}\zeta$ (17)
$\frac{\partial n}{\partial t}+\{\phi,n\}+\kappa\frac{\partial\phi}{\partial y}$ $=\alpha(\tilde{\phi}-\tilde{n})-\mu\nabla^{4}n$ (18)

### v3

@@ 2.3 Incorporating simultaneous learning of data reliability in EOMC
Learning piece-wise low-dimensional linear manifold approximations with EOMC implicitly deploys an assumption that all of the data instances have the same uniform probability $(1/T)$ to contribute to the loss function $\mathcal{L}^{\textrm{EOMC}}$ in (17-20). This is not a problem when this assumption is justified, and when the training data is spread more-or-less uniformly and close to the manifold. But, this becomes a problem when the new data instances $X$ (f.e., the ones that were not used in the training), come from somewhere arbitrarily in the feature space - and further away from the manifold. In the Sec.3.1.3 below, we provide an example of such a situation - resulting in projection artefacts, when training the manifold with the data close to the manifold (red crosses in Fig. 4), and then using the trained EOMC manifold projectors (33) from Theorem 1 for the data points that are far away from the manifold and from the training data (blue dots in Fig. 4B). This input data reliability issue is a very general and severe problem in contemporary AI, resulting in low robustness, "over-confidence" in provided answers, and in AI vulnerability to the so-called adversarial attacks, when subject to new input data that is sufficiently-different from the domain $\Omega$ of "admissible" data implicitly defined by the training data.
To address this issue - and following the ideas originally introduced in the Entropic Outlier Sparsification (EOS) and Entropy-Optimal Networks (EON) methods [21, 24] - we propose to incorporate in the EOMC problem formulation (17-20) an additional learnable variable $\gamma_{0}(t)=\mathbb{P}\left[X(:,t)\in\Omega\right]$, measuring a probability that a new data instance $X(:,t)$ belongs to the "admissible" data domain $\Omega$. Denoting $l(t)=-\sum_{k=1}^{K}\gamma(k,t)\left[\left(X(:,t)-\mu_{k}\right)^{\dagger}\left(\left(I+\alpha\right)-\mathcal{T}_{k}\mathcal{T}_{k}^{\dagger}\right)\left(X(:,t)-\mu_{k}\right)+\beta\textrm{log}_{K}\left(\gamma(k,t)\right)\right]$, we can now rewrite (17-20) as a maximisation of the expectation $\mathbb{E}_{\gamma_{0}}\left[l\right]$ of the loss $l$ wrt. this probability measure $\gamma_{0}$, subject to the Shannon entropy regularization terms $\beta_{0}\gamma_{0}\left(t\right)\textrm{log}_{T}\gamma_{0}\left(t\right)$, where $\beta_{0}\geq 0$. Re-arranging terms we obtain:
$\left\{\gamma_{0}^{*},\gamma^{*},\mu^{*}_{1},\mathcal{T}^{*}_{1},\dots,\mu^{*}_{K},\mathcal{T}^{*}_{K}\right\}=\arg\min\mathcal{L}_{\gamma_{0}}^{\textrm{EOMC}}$ (37)
$\mathcal{L}_{\gamma_{0}}^{\textrm{EOMC}}$ $=$ $\sum_{t=1}^{T}\gamma_{0}(t)\left[\beta_{0}\textrm{log}_{T}\gamma_{0}\left(t\right)-l(t)\right],$ (38)
@@
Shannon entropy regularization terms $\beta_{0}\gamma_{0}\left(t\right)\textrm{log}_{T}\gamma_{0}\left(t\right)$ are incorporated here with a positive sign, since for $\beta_{0}\equiv 0$, $\gamma_{0}$ solution of (37-42) is entropy-minimal, attaining only $1/0$ values - and the entropy-optimal $\gamma_{0}$-solutions can only be found when maximizing the Shannon entropy. It is also straightforward to validate that the previous formulation (17-20) becomes a special case of (17-20) when $\beta_{0}\rightarrow+\infty$ (since then $\gamma_{0}(t)\rightarrow(1/T)$, for all $t$).
Then, following the same line of argumentation as in the proof of Theorem 1 above, it is easy to show that for the fixed values of the loss $l$ (hence, when all other learning variables except of $\gamma_{0}$ are fixed), and for any $\beta_{0}>0$, there exists a unique solution $\gamma^{*}_{0}$ of the problem (37-42):
@@ 2.4 Selection of EOMC hyper-parameters $d,K,\alpha,\beta,\beta_{0}$
As will be demonstrated below on practical examples (for example, see Fig. 1), tuning the EOMC hyper-parameters $d$ and $K$ can be done iteratively. To select the optimal values for $d$ and $K$, for each of the $k=1,\dots,K$ manifolds one inspects the decay of eigenvalues $\Lambda^{(k)}_{1,1},\Lambda^{(k)}_{2,2},\dots,\Lambda^{(k)}_{D,D}$ for the local weighted covariances $\mathrm{Cov}_{k}\left(X,\gamma,\mu_{k}^{*}\right)$ in the local manifolds: as follows from the Theorem 1 above, $\sum_{i=d+1}^{D}\Lambda^{(k)}_{i,i}$ quantifies the amount of squared 2-norm loss from projecting the data, that is "assigned" to this manifold $k$, through the EOMC internal coordinates $\gamma(k,:)$. Hence, it quantifies the local quality of lossy compression, that is achieved when projecting the high-dimensional data on the local low-dimensional manifolds. Since (17-20) is an unsupervised learning problem, we can not apply most of the standard hyper-paremeter selection routines from ML and AI, like cross-validation and Bayesian hyper-paremeter tuning [25]. Instead, one can use here the hyper-parameter selection from the unsupervised regularization methods in statistics and computational science, like the $L$-curve method [26]. Alternatively, in the following examples we will adopt a nonparametric information-theoretic perspective to model selection - aiming to find the hyper-parameter combinations that lead to the models combining simplicity (measured as a high lossy compression rate, computed as a ratio between the raw data complexity and EOMC model descriptor length), and quality (measured as the relative loss of compression). In contrast to common parametric measures from information theory (like Akaike and Bayesian Information Criteria) [27] that rely on validity of parametric assumptions like Gaussianity, this non-parametric model selection procedure based on lossy compression allows a robust identification of good hyper-parameter combinations across all of the benchmarks considered below. As will also be shown on application examples below, EOMC does not require a careful and precise adjustment, since the results remain robust in the broad ranges of hyper-parameters.
@@ 3.1 Synthetic examples
  • Phase 1: select the hyper-parameters $d,K,\alpha,\beta$ (and $\beta_{0}$, if simultaneous learning of data reliability $\gamma_{0}$ with (37-42) is incorporated in EOMC [^3] Here, and in the next example described in Sec. 3.1.2, we perform the analysis setting $\gamma_{0}\equiv 1/T$ and $\beta_{0}=0$. An example of EOMC with simultaneous learning of the entropy-optimal data reliability function $\gamma_{0}$ will be provided in Sec. 3.1.3), as described in Sec. 2.4. Next, perform the EOMC analysis according to the Theorem 1 above. Note that the application of the standard linear PCA (left panel in Fig. 1) does not reveal any spectral gap in the eigenvalues of the covariance, and does not indicate a presence of any low-dimensional manifold. In contrast, applying EOMC with arbitrarily selected $K=15$ indicates a clear spectral gap after the second eigenvalues of the localized EOMC covariance matrices $\mathrm{Cov}_{k}\left(X,\gamma,\mu_{k}^{*}\right)$. This clearly indicates that the data contains non-linear manifolds with local dimensionality $d=2$.
@@
Despite of the hyper-parameter adjustment, PCA and t-SNE fail to recover the two low-dimensional manifolds from these nonstationary data (see Fig. 2A and 2B). Fig. 2C illustrates further effects of hyper-parameter selection described in Sec. 2.4: setting both $\alpha=0$ and $\beta=0$ makes (17-20) to its special case, i.e., to PCA-clustering [16, 17, 18]. It makes visible the central problem of PCA-clustering, discussed above and induced by the Lemmas 1 and 2: the kernel of the manifold projection operator is non-empty when $d<D$ (Lemma 1), and the errors orthogonal to the manifold are not visible to the method. Therefore, the approximation in this case is rather cutting through the manifold then approximating it. Setting $\alpha$ and $\beta$ to some non-zero values mitigates this problem (see Fig. 2D). Then, increasing the number $K$ of local manifolds from 15 (which was the initial guess) to 50 results in almost perfect reconstruction of both nonlinear manifolds (see Fig. 2D). Please note that obtaining this result did not require a tedious hyper-parameter tuning, results shown in Fig. 2C were obtained from only two repetitions of the EOMC pipeline described in the Fig. 1.
@@
As in the previous example, despite of the hyper-parameter tuning, both linear PCA and t-SNE fail to find these low dimensional manifolds hidden in 100 dimensions of noisy data (see Fig. 3A and 3B). Instead, t-SNE finds a lot of clusters (see Fig. 3B), that are not present in the data, representing a clear artefact of t-SNE. Fig. 3C illustrates again the problem of PCA-clustering, and Figs. 3D and 3E show the effect of the entropic regularization term in (17-20): setting $\alpha>0$ and $\beta=0$ results in "edgy" piece-wise linear approximation of the nonlinear manifold fragments (Fig. 3D), whereas setting $\alpha>0$ and $\beta>0$ allows obtaining much better and more smooth interpolations. Like in the example 1 above, obtaining the result from the Fig. 3E did not require a tedious hyper-parameter adjustment, requiring to go through the EOMC pipeline described in the Fig. 1 only once.
@@
Next, we illustrate the application of the modified EOMC (with a simultaneous learning the input data reliability $\gamma_{0}$), as described in Sec.  2.3 and eqs. (37-43). We generate a noisy 2D data (red crosses in Fig. 4A), that are normally distributed around a smooth 1D non-linear manifold following a smooth letter "S" (black line in Fig. 4A). As in the previous synthetic examples, to make the manifold learning more challenging, we add an additional 10-dimensional Gaussian noise and randomly rotate the training data in 10 dimensions. As discussed in Sec.  2.3, the EOMC formulation in (17-20) implicitly assumes that all of the training data points have the same $1/T$ contribution to the loss function - i.e., according to the uniform probability distribution prior. However, this assumption can be seriously violated when considering the data points not used in training and distributed uniformly inside of the whole 10-dimensional hypercube - and far away from the training data (blue dots in Fig. 4). As can be seen from Fig. 4B, applying the projection formula (33) from the Theorem 1 results in artefact projections that are far away of the true manifold (red circular dots in Fig. 4B).
In contrast, co-inference of data reliability function $\gamma_{0}(t)=\mathbb{P}\left[X(:,t)\in\Omega\right]$ in the modified EOMC from Sec.  2.3 and eqs. (37-43) produces a very reliable reconstruction of the true manifold: the intensity of red for circular dots in Fig. 4C is proportional to the inferred values of data reliability $\gamma_{0}(t)$. The unreliable points far away from the training data and from the EOMC-approximated manifold automatically get values of $\gamma_{0}(t)$ close to zero. As in the previous two synthetic examples, achieving this result did not require any particular hyper-parameter tuning: it was obtained for an ad hoc choice of $K=6,d=1,\alpha=5,\beta=0.3,\beta_{0}=4$. This result remains remarkably-robust, and does not change notably when changing these hyper-parameter values in quite broad ranges.
Figure 4: Illustration of the modified EOMC learning from Sec.  2.3 and eqs. (37-43): co-inference of data reliability function $\gamma_{0}$.
@@ 3.2 Analysis of data from Lorenz-63 in chaotic, strongly-chaotic and very strongly-chaotic regimes
We inspect the time series of intrinsic EONC coordinates $\gamma$ that were computed with these optimal hyper-parameter choices. $\gamma$ encode the probabilities of points to belong to different local manifolds $\left\{\mu_{k},\mathcal{T}_{k}\right\}$, $k=1,\dots,K$ at different time series instances. As can be seen from the right column of Fig. 5, $\gamma$ exhibits very persistent and metastable dynamics in time, indicating relatively-long stays in each of the local linear manifolds (with $d=3$). Please note that this metastability of $\gamma$ can not be an artefact of the EOMC data analysis: as can be seen from (17-20), the EOMC loss function does not contain any terms that would enforce the persistence or metastability on $\gamma$. As a matter of fact, the value of $\mathcal{L}^{\textrm{EOMC}}$ in (17-20) is invariant with respect to any permutation of the columns of data matrix $X$. This means that the observed metastability of $\gamma$ can only be an imprint of the underlying L96 dynamics - that appears to be best described by a metastable process switching between low-dimensional (with $d=3$) locally-linear manifolds. Next, for each of the obtained $\gamma$, we compute the Markovian transition operators $P$ that describe the time evolution of $\gamma(:,t)$, by means of the exact law of the total probability, i.e., $\gamma(:,t+\tau)=P\gamma(:,t)$, where $P_{i,j}=\mathbb{P}\left[\gamma(j,t+\tau)=1|\gamma(i,t)=1\right]$ [34]. Elements $P_{i,j}$ of the transfer operator $P$ contain the probabilities of transitions from state $j$ to state $i$ in a single time step $\tau$. As demonstrated by the Fig. 6A, operators $P$ are characterized by high probabilities of staying in the states (large diagonal entries) and low probabilities of transitions to other local manifold states (very small off-diagonal entries). Next, for each of the three transfer operators computed for each of the three L96 forcing regimes, we compute the relaxation times $\tau^{rel}(P)=\frac{\tau}{1-|\lambda_{2}|}$, where $\tau$ is the time step, and $\lambda_{2}$ is the second largest (in absolute value) eigenvalue of $P$ [34]. Relaxation times $\tau^{rel}(P)$ in Markov processes measure the predictability horizons: they quantify the time it takes for the Markov processes $P$ to forget the initial condition, and to converge towards the invariant density measure [35, 36, 34]. As can be seen from the Fig. 6A, $\tau^{rel}(P)$ gradually reduces from $0.73$, over $0.61$ and to $0.57$ when $F$ goes from $7$, over $9$, and to $12$.
@@
$\frac{\partial\zeta}{\partial t}+\{\phi,\zeta\}$ $=\alpha(\tilde{\phi}-\tilde{n})-\mu\nabla^{4}\zeta,$ (45)
$\frac{\partial n}{\partial t}+\{\phi,n\}+\kappa\frac{\partial\phi}{\partial y}$ $=\alpha(\tilde{\phi}-\tilde{n})-\mu\nabla^{4}n,$ (46)